# Christian Döding [^2] Institut für Numerische Simulation, Universität Bonn, Friedrich-Hirzebruch-Allee 7, D-53115 Bonn, Germany    Barbara Verfürth [^2] Institut für Numerische Simulation, Universität Bonn, Friedrich-Hirzebruch-Allee 7, D-53115 Bonn, Germany

### v2

@@ 5.3 Magnetic higher order corrections
where $\zeta$ is the unique $1$-periodic solution of the third cell-problem
@@
$\partial_{\tau}(\varepsilon\partial_{\tau}\mathbf{u}_{4})+\partial_{t}(\varepsilon\partial_{\tau}\mathbf{u}_{3})+\partial_{\tau}(\varepsilon\partial_{t}\mathbf{u}_{3})+\partial_{t}(\varepsilon\partial_{t}\mathbf{u}_{2})-\Delta\mathbf{u}_{2}=0.$
Integration over $(0,1)$ w.r.t. $\tau$ yields with our derived expression for $\mathbf{u}_{3}$ and $\mathbf{u}_{2}$
@@
$\varepsilon(\tau)\partial_{\tau}\mathbf{u}_{3}(x,t,\tau)-\varepsilon(0)\partial_{\tau}\mathbf{u}_{3}(x,t,0)$ $=\Big(\varepsilon(0)\xi(0)-\varepsilon(\tau)\xi(\tau)+\varepsilon_{\hom}\int_{0}^{\tau}\chi\,ds\Big)\partial_{ttt}\mathbf{u}_{0}(x,t)$
$\quad+(\varepsilon(0)\chi_{0}-\varepsilon(\tau)\chi(\tau))\partial_{tt}\overline{\mathbf{u}}_{1}(x,t)+(\varepsilon(0)-\varepsilon(\tau))\partial_{t}\overline{\mathbf{u}}_{2}(x,t).$
Dividing by $\varepsilon(\tau)$ and by $\varepsilon(0)$ yields, after integration over $(0,1)$ w.r.t. $\tau$,
$-\partial_{\tau}\mathbf{u}_{3}(x,t,0)$ $=-\theta_{0}\partial_{ttt}\mathbf{u}_{0}(x,t)+\frac{\varepsilon_{\hom}^{2}}{\varepsilon(0)}\int_{0}^{1}\varepsilon^{-1}\int_{0}^{\tau}\chi\,ds\,\partial_{ttt}\mathbf{u}_{0}(x,t)$
$\quad+\chi_{0}\partial_{tt}\overline{\mathbf{u}}_{1}(x,t)+\partial_{t}\overline{\mathbf{u}}_{2}(x,t)-\frac{\varepsilon_{\hom}}{\varepsilon(0)}\partial_{t}\overline{\mathbf{u}}_{2}(x,t).$
With $\partial_{t}\overline{\mathbf{u}}_{2}(x,0)=-\partial_{\tau}\mathbf{u}_{3}(x,0,0)-\partial_{t}\widetilde{\mathbf{u}}_{2}(x,0,0)$ and $\partial_{t}\widetilde{\mathbf{u}}_{2}(x,0,0)=-\theta_{0}\partial_{ttt}\mathbf{u}_{0}(x,0)+\chi_{0}\partial_{tt}\overline{\mathbf{u}}_{1}(x,0)$ this implies
@@ 6 Homogenized Maxwell’s equations
$\varepsilon_{\hom}\partial_{tt}\overline{\mathbf{D}}^{(2)}-\Delta\overline{\mathbf{D}}^{(2)}=\eta^{2}\varepsilon_{\mathrm{cor}}\Delta^{2}\mathbf{D}_{0}=\eta^{2}\varepsilon_{\hom}\varepsilon_{\mathrm{cor}}\Delta\partial_{tt}\mathbf{D}_{0}.$
Replacing $\mathbf{D}_{0}$ by $\overline{\mathbf{D}}^{(2)}-\eta\mathbf{D}_{1}-\eta^{2}\overline{\mathbf{D}}_{2}$, we obtain
$(\varepsilon_{\hom}-\varepsilon_{\hom}\varepsilon_{\mathrm{cor}}\Delta)\partial_{tt}\overline{\mathbf{D}}^{(2)}-\Delta\overline{\mathbf{D}}^{(2)}=\eta^{3}\varepsilon_{\hom}\varepsilon_{\mathrm{cor}}\Delta\partial_{tt}\mathbf{D}_{1}+\eta^{4}\varepsilon_{\hom}\varepsilon_{\mathrm{cor}}\Delta\partial_{tt}\overline{\mathbf{D}}_{2}.$
Dropping the $\mathcal{O}(\eta^{3})$ source terms on the right hand side, we define, in analogy of Section 4.4, the macroscopic electric quantity $\mathbf{D}^{(2)}$ as the solution of
@@
between the macroscopic electric displacement $\mathbf{D}^{(2)}$ and the associated macroscopic electric field $\mathbf{E}^{(2)}$. In the limit $\eta\rightarrow 0$, this gives $\mathbf{D}^{(2)}=\varepsilon_{\hom}\mathbf{E}^{(2)}$, which is the same local constitutive relation as that of the effective fields. However, for $\eta>0$ there is an additional contribution of $\eta^{2}\varepsilon_{\hom}\varepsilon_{\mathrm{cor}}\operatorname{\operatorname{curl}}\operatorname{\operatorname{curl}}\mathbf{E}^{(2)}$. This additional contribution is nonlocal, as it involves derivatives of the electric field. For monochromatic plane waves, this nonlocal material law produces artificial magnetism in the material by altering the effective magnetic permeability obtained from the dispersion relation; cf. [24, 27].
@@ 7 Numerical experiments
We first consider the electric displacement and compute the full wave problem (2.8) for the choice $\eta=2^{-4}T$. Here we use $N=64$ Fourier modes, step size $\tau=2^{-8}T$ and plot the full wave solution $\mathbf{u}_{\eta}$ in Figure 7.1 (top left). As expected, we see the wave propagating to the right accompanied by a smaller reflected wave moving to the left. The latter comes from the interface condition at $t=0$ described by the initial values. This reflection is induced when the permittivity of the metamaterial is switched from constant permittivity ($t<0$) to time-varying permittivity ($t>0$). Visually, no fine oscillation can be observed, but the propagation velocity is certainly governed by the time variation of the permittivity.

### v3

@@ 5.3 Magnetic higher order corrections
where $\zeta$ is the unique $1$-periodic solution of the third cell-problem, c.f. Table 1,
@@
$\partial_{\tau}(\varepsilon\partial_{\tau}\mathbf{u}_{4})+\partial_{t}(\varepsilon\partial_{\tau}\mathbf{u}_{3})+\partial_{\tau}(\varepsilon\partial_{t}\mathbf{u}_{3})+\partial_{t}(\varepsilon\partial_{t}\mathbf{u}_{2})-\Delta\mathbf{u}_{2}=0.$ (5.14)
Integration of (5.14) over $(0,1)$ w.r.t. $\tau$ yields with our derived expression for $\mathbf{u}_{3}$ and $\mathbf{u}_{2}$
@@
$\begin{split}\varepsilon(\tau)\partial_{\tau}\mathbf{u}_{3}(x,t,\tau)-\varepsilon(0)\partial_{\tau}\mathbf{u}_{3}(x,t,0)&=\Big(\varepsilon(0)\xi(0)-\varepsilon(\tau)\xi(\tau)+\varepsilon_{\hom}\int_{0}^{\tau}\chi\,ds\Big)\partial_{ttt}\mathbf{u}_{0}(x,t)\\ &\quad+(\varepsilon(0)\chi_{0}-\varepsilon(\tau)\chi(\tau))\partial_{tt}\overline{\mathbf{u}}_{1}(x,t)+(\varepsilon(0)-\varepsilon(\tau))\partial_{t}\overline{\mathbf{u}}_{2}(x,t).\end{split}$ (5.15)
Dividing (5.15) by $\varepsilon(\tau)$ and by $\varepsilon(0)$ yields, after integration over $(0,1)$ w.r.t. $\tau$,
$\begin{split}-\partial_{\tau}\mathbf{u}_{3}(x,t,0)&=-\theta_{0}\partial_{ttt}\mathbf{u}_{0}(x,t)+\frac{\varepsilon_{\hom}^{2}}{\varepsilon(0)}\int_{0}^{1}\varepsilon^{-1}\int_{0}^{\tau}\chi\,ds\,\partial_{ttt}\mathbf{u}_{0}(x,t)\\ &\quad+\chi_{0}\partial_{tt}\overline{\mathbf{u}}_{1}(x,t)+\partial_{t}\overline{\mathbf{u}}_{2}(x,t)-\frac{\varepsilon_{\hom}}{\varepsilon(0)}\partial_{t}\overline{\mathbf{u}}_{2}(x,t).\end{split}$ (5.16)
With $\partial_{t}\overline{\mathbf{u}}_{2}(x,0)=-\partial_{\tau}\mathbf{u}_{3}(x,0,0)-\partial_{t}\widetilde{\mathbf{u}}_{2}(x,0,0)$ and $\partial_{t}\widetilde{\mathbf{u}}_{2}(x,0,0)=-\theta_{0}\partial_{ttt}\mathbf{u}_{0}(x,0)+\chi_{0}\partial_{tt}\overline{\mathbf{u}}_{1}(x,0)$ equation (5.16) implies
@@ 5.4 Pure macroscopic description of the magnetic field
We close this section with a discussion of our findings, focusing in particular on the comparison between the electric and magnetic cases. At the effective level, both fields satisfy the homogenized wave equation with the effective coefficient $\varepsilon_{\hom}$ and the respective pure initial values $\mathbf{v}^{\mathrm{e}}_{0},\mathbf{v}^{\mathrm{e}}_{1}$ for the electric field, and $\mathbf{v}^{\mathrm{m}}_{0},\mathbf{v}^{\mathrm{m}}_{1}$ for the magnetic field. The first notable differences appear in the first-order corrections: in the electric case, no microscopic dynamics are observed, whereas in the magnetic case, microscopic dynamics arise through the solution $\chi$ of the first cell problem (5.2). At the macroscopic level, both first-order corrections again satisfy the homogenized wave equation with coefficient $\varepsilon_{\hom}$, but now with modified initial values induced by the interface conditions (see Conclusions 2 and 5). Finally, for the second-order corrections in both fields, microscopic dynamics are induced via solutions of higher-order cell problems. At the macroscopic level, we obtain, up to a sign, the same homogenized PDEs for the second-order components, now featuring higher-order (nonlocal) source terms and appropriately calculated initial values.
@@ 6 Homogenized Maxwell’s equations
$\varepsilon_{\hom}\partial_{tt}\overline{\mathbf{D}}^{(2)}-\Delta\overline{\mathbf{D}}^{(2)}=\eta^{2}\varepsilon_{\mathrm{cor}}\Delta^{2}\mathbf{D}_{0}=\eta^{2}\varepsilon_{\hom}\varepsilon_{\mathrm{cor}}\Delta\partial_{tt}\mathbf{D}_{0}.$ (6.2)
Replacing $\mathbf{D}_{0}$ by $\overline{\mathbf{D}}^{(2)}-\eta\mathbf{D}_{1}-\eta^{2}\overline{\mathbf{D}}_{2}$ in (6.2), we obtain
$(\varepsilon_{\hom}-\varepsilon_{\hom}\varepsilon_{\mathrm{cor}}\Delta)\partial_{tt}\overline{\mathbf{D}}^{(2)}-\Delta\overline{\mathbf{D}}^{(2)}=\eta^{3}\varepsilon_{\hom}\varepsilon_{\mathrm{cor}}\Delta\partial_{tt}\mathbf{D}_{1}+\eta^{4}\varepsilon_{\hom}\varepsilon_{\mathrm{cor}}\Delta\partial_{tt}\overline{\mathbf{D}}_{2}.$ (6.3)
Dropping the $\mathcal{O}(\eta^{3})$ source terms on the right hand side of (6.3), we define, in analogy of Section 4.4, the macroscopic electric quantity $\mathbf{D}^{(2)}$ as the solution of
@@
between the macroscopic electric displacement $\mathbf{D}^{(2)}$ and the associated macroscopic electric field $\mathbf{E}^{(2)}$. In the limit $\eta\rightarrow 0$, this gives $\mathbf{D}^{(2)}=\varepsilon_{\hom}\mathbf{E}^{(2)}$, which is the same local constitutive relation as that of the effective fields. However, for $\eta>0$ there is an additional contribution of $\eta^{2}\varepsilon_{\hom}\varepsilon_{\mathrm{cor}}\operatorname{\operatorname{curl}}\operatorname{\operatorname{curl}}\mathbf{E}^{(2)}$. This additional contribution is nonlocal, as it involves derivatives of the electric field. For monochromatic plane waves, this nonlocal material law produces artificial magnetism in the material by altering the effective magnetic permeability obtained from the dispersion relation; cf. [29, 32]. As demonstrated in the numerical simulations reported in [29], incorporating the nonlocal term (artificial magnetism) into the model of light propagation can significantly improve the accuracy of the approximation in certain parameter regimes, compared to a purely local description. In the next section, we present analogous numerical observations in particular in the regime with $\omega_{0}\ll\omega_{m}$, where we recall that $\omega_{m}=\eta^{-1}$ and $\omega_{0}$ denotes the spatial modulation frequency. For a quantitative analysis of the relevant parameter regimes, we refer to the experimental study in [29].
@@ 7 Numerical experiments
We first consider the electric displacement and compute the full wave problem (2.8) for the choice $\eta=2^{-4}T$. Here we use $N=64$ Fourier modes, step size $\tau=2^{-8}T$ and plot the full wave solution $\mathbf{u}_{\eta}$ in Figure 7.1 (top left). As expected, we observe the wave propagating to the right, accompanied by a another reflected wave traveling to the left. This second component can be identified as a temporal reflection phenomenon, which is well-known to arise at the temporal interface $t=0$, see e.g. [18]. More specifically, it corresponds to a phase-conjugated component with negative frequency, generated when the metamaterial switches from a constant permittivity for $t<0$ to a time-varying permittivity for $t>0$. The temporal interface conditions thus split the wave into a forward-propagating (time-refracted) part and a backward-propagating (phase-conjugated) part. While no fine-scale oscillations are visually apparent, the propagation speed is clearly governed by the temporal modulation of the permittivity.
@@
8 Conclusion
In this work, we have developed a systematic higher-order homogenization framework for electromagnetic wave propagation in temporally modulated metamaterials with spatially homogeneous, time-dependent permittivity. Based on a formal two-scale asymptotic expansion, we derived the effective equations together with corrections up to second order, thereby capturing both local and nonlocal material laws. Our analysis reveals structural differences between the electric and magnetic field formulations and demonstrates how the position of the time-dependent coefficients determines the form of the resulting homogenized equations. A key contribution of this study is the careful derivation of consistent initial data and temporal interface conditions, which are intrinsic to the time-dependent setting and cannot be obtained through a straightforward space–time analogy. Our numerically verified formal results unify and extend previous approaches, providing a consistent mathematical framework for the effective modeling of time-varying metamaterials and offering a methodology that can be transferred and further developed for related wave systems, including acoustics, elasticity, and anisotropic metamaterials.
@@ Acknowledgement
This work is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 258734477 – SFB 1173 and under Germany’s Excellence Strategy – EXC-2047/1 – 390685813.